### v2

@@ 1 Introduction
Essentially, both the TCR CDR3 and antigens can be expressed as a sequence of amino acids, exhibiting striking similarity to human natural language. Large language models have gained increasing attention in the field of protein modeling in recent years [26, 27, 28], and yield to new insights into the understanding for structures and functions of proteins. Most protein language models are built upon the Transformer encoder[29], which can encode protein sequences or structures into predefined-length latent representations. The representations have been proven to boost the performance of downstream tasks related to proteins, such as the prediction of structures, functions, and protein-drug interactions. The masked language models (MLM), such as ESM1b [30], ESM-1v [31], ProtFlash [32] and ProtTrans [33], aim to reconstruct masked tokens based on surrounding sequences. Their success has inspired to make advantage of MLM to accommodate the diversity of TCR sequences. For example, TULIP [34] exploits the Transformer-based unsupervised language model to predict the interactions between peptides and TCRs. TAPIR [35] is also inspired by language models and involves a two-tower architecture composed of independent CNN-based encoders for TCR and target sequences. ProtLM.TCR [36] is a RoBERTa-style Transformer-based language model designed to predict binding between TCR and HLA class I epitope sequences.
@@ 2.3 Evaluation on external and COVID-19 test dataset
We further fine-tuned four other protein language models pretrained on generic protein sequences set (e.g. UniRef50 and UniRef90) to predict pTCR bindings. These language models include proteinBERT [38], ProtFlash [32], ProtTrans [33] and ESM2 [39]. We compared their parameter sizes (Fig. 4b, $\log_{2}$ scale), and found that ProTrans had the largest parameter size (3 billion), while tcrLM-S and ProteinBERT had much smaller parameter sizes. We used tcrLM-M in performance evaluation experiments, because it had slightly more parameters than ProteinBERT but much fewer than the other protein language models.
@@
For pTCR binding prediction, we fed the TCR and antigen embeddings derived from these pretrained language models into the projection head. During fine-tuning stage, the parameters of the pretrained encoders were kept frozen. To evaluate the degree of fitness to the pTCR training samples, we plotted their loss curves over training epochs (Fig. 4c). The results demonstrated that all models converged, with ProTrans achieving the lowest loss value, followed by tcrLM and ProtFlash. Next, we evaluated their performance on the independent, external and COVID-19 test sets, respectively. The results showed that tcrLM achieved superior performance on most metrics across all three test sets (Fig. 4d). Although ProTrans surpassed tcrLM on certain performance metrics, this came at the cost of a parameter size approximately 40 times larger than that of tcrLM. Other models exhibited inferior performance compared to tcrLM, underscoring tcrLM’s ability to outperform previously published protein language models in TCR-specific tasks.
@@ 2.5 Predicted TCR-antigen bindings indicates immunotherapy outcomes
To further explore the predictive capacity of tcrLM, we conducted an exploratory analysis of the predicted pTCR binding scores on a cohort of patients with advanced melanoma. This cohort consists of 29 patients who underwent treatment by immune checkpoint inhibitors. Based on the TCR-seq and genomic sequencing data, we took each missense mutation as an anchor, and generated all possible 9-mer peptides covering this anchor. After extracting CDR3 sequences from the TCR-seq data, we created all conceivable peptide-CDR3 pairs for each patient, yielding a total of 81,851,486 pTCR candidates.
We scored these candidate pairs using fine-tuned tcrLM, and selected the top 25,000 highest-scored pTCRs for each patient. By stratifying the patients into benefit, non-benefit, and long-term survival groups, we observed that patients in the long-term survival group had higher scored pTCRs compared to the other groups (Fig. 5a). Subsequently, we categorized the patients into four categories according to RECIST criteria: Complete Remission (CR), Partial Remission (PR), Stable Disease (SD), and Progressive Disease (PD) groups, and found the predicted bindings scores of four groups differed significantly from each other (Fig. 5b). The analysis of variance (ANOVA) verified the statistical differences among these groups ($F$-test, p-value¡0.01). This findings validated the correlation between higher pTCR predicted scores and better immunotherapy outcomes, offering a new perspective and a possible clinical indicator for tumor immunotherapy.
@@
Figure 5: Correlation between predicted pTCR scores and immunotherapy outcomes in a melanoma cohort. (a-b) Violin and boxplots of predicted pTCR binding scores regarding the different immunotherapy response groups. (c-d) Overall survival and progression free survival curves between stratified patient groups with high- and low-confidence pTCR bindings.
@@
Figure 6: tcrLM captured the biochemical properties and positional preference of amino acids within TCR sequences. (a) tSNE plot of the high-dimensional embeddings of 20 amino acids obtained from the stacked encoder. (b) Heatmaps of the computed cosine similarities using the BLOSUM62 substitution matrix and learned representations of of 20 amino acids by tcrLM. (d) Heatmap generated by cumulative attention weights regarding each position for 5-21mer TCR sequences. (d)
@@ 4 Materials and Methods
4.1 Data source and preprocessing
@@
4.3 Predicting pTCR binding
The pre-trained encoder is used for the prediction task of pTCR binding. As shown in Fig. 1(b), the antigen sequences and TCR sequences are padded to the maximum length of 15 and 20, respectively. After passing through an embedding layer, these two sequences are taken as input of the pre-trained encoder, and transformed into matrices of dimensions 15x512 and 20x512. Subsequently, these two matrices are concatenated along the feature dimension to form a 35x512 matrix, which is then flattened into a 17,920-dimensional vector. This vector passes through a fully-connected layer and is projected to a 2-dimensional vector, followed by softmax layer to obtain the final predicted probabilities. During the fine-tuning process, the pre-trained encoder is frozen. For pTCR binding prediction task, we use the cross-entropy as the loss function:
@@ 4.3 Predicting pTCR binding
4.4 Virtual adversarial training
@@
$D$ represents the function that measures the divergence between two distributions, $p(y|x)$ denotes the probability of the model predicting label $y$ given input $x$, $r_{\mathrm{vadv}}$ is a virtual adversarial perturbation regarding the input sample $x$. This perturbation strives to maximize the divergence between $p(y|x_{*},\hat{\theta})$ and $p(y|x+r)$ by choosing the direction of gradient ascent.

### v3

@@ 1 Introduction
Essentially, both the TCR CDR3 and antigens can be expressed as a sequence of amino acids, exhibiting striking similarity to human natural language. Large language models have gained increasing attention in the field of protein modeling in recent years [26, 27, 28], and yield to new insights into the understanding for structures and functions of proteins. Most protein language models are built upon the Transformer encoder[29], which can encode protein sequences or structures into predefined-length latent representations. The representations have been proven to boost the performance of downstream tasks related to proteins, such as the prediction of structures, functions, and protein-drug interactions. The masked language models (MLM), such as CVC [30], TCR-BERT [31], ESM1b [32], ESM-1v [33], ProtFlash [34] and ProtTrans [35], aim to reconstruct masked tokens based on surrounding sequences. Their success has inspired to make advantage of MLM to accommodate the diversity of TCR sequences. For example, TULIP [36] exploits the Transformer-based unsupervised language model to predict the interactions between peptides and TCRs. TAPIR [37] is also inspired by language models and involves a two-tower architecture composed of independent CNN-based encoders for TCR and target sequences. ProtLM.TCR [38] is a RoBERTa-style Transformer-based language model designed to predict binding between TCR and HLA class I epitope sequences. CVC [30] uses a lightly modified BERT architecture with tweaked pre-training objectives trained on TCR CDR3 sequences. TCR-BERT [31] adopts self-supervised learning to learn latent representations of TCR CDR3 sequences, thereby enhancing the performance of downstream tasks, such as TCR-antigen binding prediction, sequence clustering, and engineered TCR design.
@@ 2.3 Evaluation on external and COVID-19 test dataset
We further fine-tuned four other protein language models pretrained on generic protein sequences set (e.g. UniRef50 and UniRef90) to predict pTCR bindings. These language models include proteinBERT [40], ProtFlash [34], ProstT5 [41], ESM2 [42], CVC [30] and TCR-BERT [31]. We compared their parameter sizes (Fig. 4b, $\log_{2}$ scale), and found that ProstT5 had the largest parameter size (3 billion), while tcrLM-S and ProteinBERT had much smaller parameter sizes. We used tcrLM-M in performance evaluation experiments, because it had slightly more parameters than ProteinBERT but much fewer than the other protein language models.
@@
For pTCR binding prediction, we fed the TCR and antigen embeddings derived from these pretrained language models into the projection head. During fine-tuning stage, the parameters of the pretrained encoders were kept frozen. To evaluate the degree of fitness to the pTCR training samples, we plotted their loss curves over training epochs (Fig. 4c). The results demonstrated that all models converged, with CVC achieving the lowest loss value, followed by TCR-BERT and ProstT5. Next, we evaluated their performance on the independent, external and COVID-19 test sets, respectively. The results showed that tcrLM achieved superior performance and generalization capability on most metrics across all three test sets (Fig. 4d). Although ProstT5 surpassed tcrLM on certain performance metrics, this came at the cost of a parameter size approximately 40 times larger than that of tcrLM. CVC and TCR-BERT demonstrate comparable performance to tcrLM on independent and external test sets. However, their performance on the COVID-19 test set falls short of tcrLM, indicating slightly inferior generalization capabilities. Other models exhibited inferior performance compared to tcrLM, underscoring tcrLM’s ability to outperform previously published protein language models in TCR-specific tasks.
@@ 2.5 Predicted TCR-antigen bindings indicates immunotherapy outcomes
To further explore the predictive capacity of tcrLM, we conducted an exploratory analysis of the predicted pTCR binding scores on a cohort of patients with advanced melanoma [43]. This cohort consists of 29 patients who underwent treatment by immune checkpoint inhibitors. Based on the TCR-seq and genomic sequencing data, we took each missense mutation as an anchor, and generated all possible 9-mer peptides covering this anchor. After extracting CDR3 sequences from the TCR-seq data, we created all conceivable peptide-CDR3 pairs for each patient, yielding a total of 81,851,486 pTCR candidates.
We scored these candidate pairs using fine-tuned tcrLM, and selected the top 25,000 highest-scored pTCRs for each patient. By stratifying the patients into benefit, non-benefit, and long-term survival groups, we observed that patients in the long-term survival group had higher scored pTCRs compared to the other groups (Fig. 5a). Subsequently, we categorized the patients into four categories according to RECIST criteria: Complete Remission (CR, nn=2), Partial Remission (PR, nn=5), Stable Disease (SD, nn=9), and Progressive Disease (PD, nn=12) groups, and found the predicted bindings scores of four groups differed significantly from each other (Fig. 5b). The analysis of variance (ANOVA) verified the statistical differences among these groups ($F$-test, p-value¡0.01). This findings validated the correlation between higher pTCR predicted scores and better immunotherapy outcomes, offering a new perspective and a possible clinical indicator for tumor immunotherapy.
@@
Figure 5: Correlation between predicted pTCR scores and immunotherapy outcomes in a melanoma cohort. (a-b) Violin and boxplots of predicted pTCR binding scores regarding the different immunotherapy response groups (p-value¡0.01, $F$-tests). (c-d) Overall survival and progression free survival curves between stratified patient groups with high- and low-confidence pTCR bindings.
@@
Figure 6: tcrLM captured the biochemical properties and positional preference of amino acids within TCR sequences. (a) tSNE plot of the high-dimensional embeddings of 20 amino acids obtained from the stacked encoder. (b) Heatmaps of the computed cosine similarities using the BLOSUM62 substitution matrix and learned representations of of 20 amino acids by tcrLM. (c) Heatmap generated by cumulative attention weights regarding each position for 5-21mer TCR sequences. (d) Sequence logo of the CDR3 sequences where the relative importance of amino acids at each position is quantified by the cumulative attention weights.
@@ 4 Materials and Methods
Data source and preprocessing
@@
For further performance validation, we establish an external test set comprising 63,324 pTCR bindings across 890 unique peptides and 53,439 CDR3 sequences collected from numbers of publications. In addition, a large-scale test set are sourced from ImmuneCODE database [39], consisting of 520,000 pTCR bindings between COVID-19 virus-derived antigens and human TCRs.
@@
Predicting pTCR binding
The pre-trained encoder is subsequently used for the prediction task of pTCR binding. As shown in Fig. 1(b), the antigen sequences and TCR sequences are padded to the maximum length of 15 and 20, respectively. After passing through an embedding layer, these two sequences are taken as input of the pre-trained encoder, and transformed into matrices of dimensions 15x512 and 20x512. Subsequently, these two matrices are concatenated along the feature dimension to form a 35x512 matrix, which is then flattened into a 17,920-dimensional vector. This vector passes through a fully-connected layer and is projected to a 2-dimensional vector, followed by softmax layer to obtain the final predicted probabilities. During the fine-tuning process, the pre-trained encoder is frozen. For pTCR binding prediction task, we use the cross-entropy as the loss function:
@@
Virtual adversarial training
@@
$\text{ where }r_{\mathrm{vadv}}$ $=\arg\max_{r;\|r\|\leq\epsilon}D\left[p(y|x_{*},\hat{\theta}),p(y|x+r)\right],$